### 06 Apr 2026 v1

@@ 4.2 Two-phase path-following framework
From remarks (ii) and (iii), when Algorithm 2 terminates, an approximate KKT solution $(x^{(k)},s^{(k)},\lambda^{(k)})$ to the original SCP problem (1) is obtained, satisfying
$\mathcal{A}x^{(k)}=b,\;\mathcal{A}^{*}\lambda^{(k)}+s^{(k)}=c,\;\|\Phi_{\rho}(x^{(k)},s^{(k)};\mu^{(k)})\|\leq 2\sqrt{\frac{\varepsilon}{\rho}}\kappa.$ (94)
4.3 Linear system in the algorithm
During practical computations, the primal and dual constraints are satisfied at all iterates (see (91)). Thus, the Newton systems (85) and (45) arising in Algorithms 1 and 2 can be written in the following unified form:
$\begin{pmatrix}\mathcal{A}&0&0\\ 0&\mathcal{I}_{\mathbb{E}}&\mathcal{A}^{*}\\ \mathcal{H}^{-1}\mathcal{W}&\rho^{-1}\mathcal{H}^{-1}&0\end{pmatrix}\begin{pmatrix}\Delta x\\ \Delta s\\ \Delta\lambda\end{pmatrix}=\begin{pmatrix}0\\ 0\\ r\end{pmatrix}$, (95)
@@ 4.3 Linear system in the algorithm
Directly solving the full Newton system (95) is computationally expensive, particularly for large-scale problems where the dimensions of primal and dual variables $x$ and $s$ are significantly larger than the number of constraints. To reduce the computational cost, we eliminate $\Delta x$ and $\Delta s$ from (95), which leads to the following reduced system:
$\mathcal{A}\mathcal{W}^{-1}\mathcal{A}^{*}\Delta\lambda$ $=-\rho\mathcal{A}\mathcal{W}^{-1}\mathcal{H}r,$ (96a)
$\Delta s$ $=-\mathcal{A}^{*}\Delta\lambda,$ (96b)
$\Delta x$ $=\mathcal{W}^{-1}(\mathcal{H}r-\rho^{-1}\Delta s).$ (96c)
For both IPMs and classical SNMs, the dominant computational cost typically comes from forming and factorizing the Schur complement $\mathcal{A}\mathcal{D}\mathcal{A}^{*}$. In our method, the matrix takes the form $\mathcal{D}=\mathcal{W}^{-1}$, where $\mathcal{W}$ is an iterate-dependent matrix determined by the barrier function in PFSNM. Accordingly, improving the efficiency of constructing $\mathcal{A}\mathcal{D}\mathcal{A}^{*}$ is crucial for large-scale performance. To this end, Proposition 3 provides closed-form expressions for $\mathcal{D}$ (and hence for $\mathcal{A}\mathcal{D}\mathcal{A}^{*}$) in PFSNM for the three most common symmetric cones.
@@
Let $\mathbb{K}$ be one of the symmetric cones listed below, and let $e$ denote the corresponding Jordan identity element. Then the corresponding Schur complement admits the following explicit representations.
@@
    Let $\mathbb{K}=\mathbb{Q}^{n+1}$. Then, for any $z\in\operatorname{int}\,(\mathbb{Q}^{n+1})$, $\phi(z)=-\frac{1}{2}\ln\big(z_{0}^{2}-\|\bar{z}\|^{2}\big),$ and
@@
By definition, ${\mathcal{W}}=\dfrac{\mu}{\rho}D^{2}\phi(z_{\rho})$. The results follow directly from the explicit formulas (see (Vieira, 2007, Proposition 2.6.1)) of $D^{2}\phi$.
The importance of having explicit formulas has already been observed in the SDP literature. The SNM in Chen and Tseng (2003), based on the smoothing FB function, has a Schur-complement formation cost comparable to that of the most expensive AHO direction in IPMs. If the smoothing CHKS function is used instead, the formation cost becomes cheaper than that of AHO, but remains more expensive than that of the NT direction. In our earlier work Zhang et al. (2024), we showed that once explicit formulas of the Schur complement are available, the formation cost can be reduced to the same order as that of the NT direction. Compared with Zhang et al. (2024), the work in this paper leverages self-concordant properties to simplify the derivation of this explicit Schur complement formation, resulting in a more direct construction.
For SOCP, the advantage of explicit Schur complement formation is even more significant. As shown in case (ii) of Proposition 3, the Schur complement takes the form of a scaled matrix combined with two rank-one updates. However, the vector $u:=\mathcal{A}z_{\rho}$ appearing in these rank-one terms is typically dense, which causes the full Schur complement to be dense as well. Figure 2 illustrates this structure and shows how these components combine to yield a fully dense matrix. This density poses a major computational challenge in large-scale settings. Even accelerating the evaluation of $\mathcal{D}$ as described in Fukushima et al. (2002) does not solve the problem, because the Schur complement remains dense in any case. The explicit representation in Proposition 3 avoids forming this dense matrix entirely. The terms $\mathcal{A}\mathcal{A}^{*}$ and $\mathcal{A}e$ depend only on the problem data and can be precomputed once. Each subsequent iteration then requires only one matrix-vector product $u=\mathcal{A}z_{\rho}$, along with simple low-rank updates and the scalar $\det(z_{\rho})$. With this structure, one can apply the product-form Cholesky factorization approach in Alizadeh and Goldfarb (2003) or the expanded sparse representation technique in Zhang et al. (2026). Both approaches exploit the low-rank structure and avoid forming a dense Schur complement.
@@ 5 Complexity analysis
In this section, we establish a polynomial iteration-complexity bound for PFSNM for SCP. The analysis is divided into two parts. We first derive a complexity bound for the first phase. Starting from the point produced by this phase, Algorithm 2 then attains a complexity bound of order $\mathcal{O}(\sqrt{\nu}\ln(1/\varepsilon))$, matching the classical short-step interior-point complexity Vavasis and Ye (1996).
The complexity bound for the first phase is presented in the following theorem.
@@
Suppose that $(\hat{x}^{(0,0)},s^{(0,0)})\in K(\theta_{\rho})$ and that $t^{(0)}$ in Algorithm 1 satisfies (87). Then Algorithm 1 requires at most
$\mathcal{O}\left(\ln\left(1+\frac{t^{(0)}\Theta_{1}(\theta_{\rho}(w^{(0,0)};\mu^{(0)}))}{\kappa}\right)\right)$ (97)
iterations to attain a starting point $v^{(0)}\in\mathcal{N}(\kappa,\mu^{(0)},\rho)$, where $\Theta_{1}(\cdot)$ is a properly chosen universal positive continuous and nondecreasing function on $\mathbb{R}_{+}$.
@@
Since $\mu$ remains fixed in the first phase, write $\eta_{\rho}(w)$ and $\eta_{t,\rho}(w)$ in place of $\eta_{\rho}(w;\mu^{(0)})$ and $\eta_{t,\rho}(w;\mu^{(0)})$, respectively. Define the merit function associated with $\eta_{t,\rho}$ by
$\delta_{t,\rho}(w):=\|(D^{2}_{ww}\eta_{\rho}(w))^{-1}\nabla_{w}\eta_{t,\rho}(w)\|_{\eta_{t,\rho},w}=\|(D^{2}_{ww}\eta_{\rho}(w))^{-1}\nabla_{w}\eta_{t,\rho}(w)\|_{\eta_{\rho},w}.$ (98)
@@
$\delta_{t^{(j+1)},\rho}(w^{(0,j)})-\delta_{t^{(j)},\rho}(w^{(0,j)})$ (100)
$=$ $\ \|(D^{2}_{ww}\eta_{\rho}(w^{(0,j)}))^{-1}\nabla_{w}\eta_{t^{(j+1)},\rho}(w^{(0,j)})\|_{\eta_{\rho,w}}$
$\qquad\qquad\qquad-\|(D^{2}_{ww}\eta_{\rho}(w^{(0,j)}))^{-1}\nabla_{w}\eta_{t^{(j)},\rho}(w^{(0,j)})\|_{\eta_{\rho},w}$
$\overset{(i)}{\leq}$ $\ \alpha^{(j)}t^{(j)}\|(D^{2}_{ww}\eta_{\rho}(w^{(0,j)}))^{-1}\nabla_{w}\eta_{\rho}(w^{(0,0)})\|_{\eta_{\rho},w^{(0,j)}}$
$\overset{(ii)}{\leq}$ $\ \frac{\kappa}{4}.$
Here, $(i)$ follows from the identity for $\nabla_{w}\eta_{t,\rho}(w)$ and that $\|\cdot\|_{\eta_{\rho},w^{(0,j)}}$ defines a norm. Inequality $(ii)$ follows from (88). Thus,
@@
$\delta_{t^{(j+1)},\rho}(w^{(0,j+1)})\leq\left(\frac{\delta_{t^{(j+1)},\rho}(w^{(0,j)})}{1-\delta_{t^{(j+1)},\rho}(w^{(0,j)})}\right)^{2}\leq\frac{\kappa}{2},$ (102)
@@
It follows from (Nemirovski, 1999, Lemma 8.3(a)), there exists a universal positive continuous and nondecreasing function $\Theta_{1}$ such that
$\|(D^{2}_{ww}\eta_{\rho}(w^{(0,j)}))^{-1}\nabla_{w}\eta_{\rho}(w^{(0,0)})\|_{\eta_{\rho},w^{(0,j)}}\leq\Theta_{1}(\theta_{\rho}(w^{(0,0)};\mu^{(0)})).$ (104)
Consequently, by (88),
$\underline{\alpha}:=\min\left\{1,\frac{\kappa}{4t^{(0)}\Theta_{1}(\theta_{\rho}(w^{(0,0)};\mu^{(0)}))}\right\}\leq\alpha^{(j)}\leq 1.$ (105)
We now consider two different cases associated with the value of $\underline{\alpha}$.
If $\underline{\alpha}=1$, then $\alpha^{(0)}=1$ and $t^{(1)}=0$. Since $\delta_{t^{(j)},\rho}(w^{(0,j)})\leq\frac{\kappa}{2}$ at each iteration $j$, we have
$\delta_{\rho}(w^{(0,1)})=\delta_{t^{(1)},\rho}(w^{(0,1)})\leq\frac{\kappa}{2}.$ (106)
Together with (90), this implies that one additional iteration leads to $v^{(0)}\in\mathcal{N}(\kappa,\mu^{(0)},\rho)$. Suppose that $\underline{\alpha}<1$. Then,
$t^{(j)}\leq(1-\underline{\alpha})^{j}t^{(0)}.$ (107)
Combining (104) and (107) gives
$\delta_{\rho}(w^{(0,j)};\mu^{(0)})$ (108)
$=$ $\ \|(D^{2}_{ww}\eta_{\rho}(w^{(0,j)}))^{-1}\nabla_{w}\eta_{\rho}(w^{(0,j)})\|_{\eta_{\rho},w^{(0,j)}}$
$\leq$ $\ \delta_{t^{(j)},\rho}(w^{(0,j)})+t^{(j)}\|(D^{2}_{ww}\eta_{\rho}(w^{(0,j)}))^{-1}\nabla_{w}\eta_{\rho}(w^{(0,0)})\|_{\eta_{\rho},w^{(0,j)}}$
$\leq$ $\ \frac{\kappa}{2}+(1-\underline{\alpha})^{j}t^{(0)}\Theta_{1}(\theta_{\rho}(w^{(0,0)};\mu^{(0)})).$
Using (90), it requires at most $\mathcal{O}\left(\ln\left(1+\frac{t^{(0)}\Theta_{1}(\theta_{\rho}(w^{(0,0)};\mu^{(0)}))}{\kappa}\right)\right)$ iterations to attain the starting point $v^{(0)}\in\mathcal{N}(\kappa,\mu^{(0)},\rho)$. This completes the proof.
Having established the bound for the first-phase of PFSNM, we now turn to analyze the complexity bound of Algorithm 2. As a preliminary step, Theorems 5.2 and 5.3 quantify the effect of updates in the barrier parameter $\mu$ on both $\nabla_{w}\eta_{\rho}(\hat{x},s;\mu)$ and $S_{\eta_{\rho}}(\hat{x},s;\mu)$. For notational simplicity, we omit the arguments $(\hat{x},s;\mu)$ and write, for example, $\nabla_{w}\eta_{\rho}:=\nabla_{w}\eta_{\rho}(\hat{x},s;\mu)$ whenever first- or second-order derivatives of $\eta_{\rho}$ with respect to $w=(\hat{x},s)$ are mentioned. All derivatives with respect to $\mu$ are denoted by a prime.
@@
For any $\mu>0$, $\rho\geq 1$, any direction $h=(h_{\hat{x}},h_{s})\in\hat{\mathbb{E}}\times\mathbb{E}$, and any point $(\hat{x},s)\in\hat{\mathbb{E}}\times\mathbb{E}$,
$\left|\langle h,\nabla_{w}\eta_{\rho}^{\prime}(\hat{x},s;\mu)\rangle\right|\leq\sqrt{\dfrac{2\nu}{\mu}}\sqrt{S_{\eta_{\rho}}(\hat{x},s;\mu)[h,h]}.$ (109)
@@
$\nabla_{w}\eta_{\rho}^{\prime}(\hat{x},s;\mu)=\begin{pmatrix}-\mathcal{B}^{*}y^{\prime}_{\rho}\\ z^{\prime}_{\rho}\end{pmatrix}=\begin{pmatrix}\mathcal{B}^{*}\mathcal{H}^{-1}\nabla\phi(z_{\rho})\\ -\rho^{-1}\mathcal{H}^{-1}\nabla\phi(z_{\rho})\end{pmatrix}$. (110)
@@
$\langle\nabla_{w}\eta_{\rho}^{\prime}(\hat{x},s;\mu),h\rangle=\langle\mathcal{B}h_{\hat{x}},\mathcal{H}^{-1}\nabla\phi(z_{\rho})\rangle-\rho^{-1}\langle h_{s},\mathcal{H}^{-1}\nabla\phi(z_{\rho})\rangle.$ (111)
@@
$|\langle\nabla_{w}\eta^{\prime}_{\rho},h\rangle|$ $\leq\left|\langle h_{x},\mathcal{H}^{-1}\nabla\phi(z_{\rho})\rangle\right|+\rho^{-1}\left|\langle h_{s},\mathcal{H}^{-1}\nabla\phi(z_{\rho})\rangle\right|$ (112)
@@
$\qquad\qquad\qquad+\rho^{-1}\sqrt{\langle h_{s},\mathcal{H}^{-1}h_{s}\rangle}\sqrt{\langle\nabla\phi(z_{\rho}),\mathcal{H}^{-1}\nabla\phi(z_{\rho})\rangle}$
$\overset{(i)}{\leq}\sqrt{\dfrac{\langle h_{\hat{x}},D_{\hat{x}\hat{x}}^{2}\eta_{\rho}h_{\hat{x}}\rangle}{\rho}}\sqrt{\dfrac{\rho}{\mu}\langle\nabla\phi(z_{\rho}),(D^{2}\phi(z_{\rho}))^{-1}\nabla\phi(z_{\rho})\rangle}$
$\qquad\qquad\qquad+\sqrt{-\dfrac{\langle h_{s},D_{ss}^{2}\eta_{\rho}h_{s}\rangle}{\rho}}\sqrt{\dfrac{\rho}{\mu}\langle\nabla\phi(z_{\rho}),(D^{2}\phi(z_{\rho}))^{-1}\nabla\phi(z_{\rho})\rangle}$
$\overset{(ii)}{=}\sqrt{\dfrac{\nu}{\mu}}\left(\sqrt{\langle h_{\hat{x}},D_{\hat{x}\hat{x}}^{2}\eta_{\rho}h_{\hat{x}}\rangle}+\sqrt{-\langle h_{s},D_{ss}^{2}\eta_{\rho}h_{s}\rangle}\right)$
@@
Here, $(i)$ follows from Corollary 1, $\mathcal{H}^{-1}\prec\mathcal{I}_{\mathbb{E}}$, and $\mathcal{H}^{-1}\prec\mathcal{W}^{-1}$. The identity $(ii)$ follows from (11).
@@
For any $\mu>0$, $\rho\geq 1$, any direction $h=(h_{\hat{x}},h_{s})\in\hat{\mathbb{E}}\times\mathbb{E}$, and any point $(\hat{x},s)\in\hat{\mathbb{E}}\times\mathbb{E}$,
$\left|S_{\eta_{\rho}}^{\prime}(\hat{x},s;\mu)[h,h]\right|\leq\frac{\rho(1+{2\sqrt{\nu}})}{\mu}S_{\eta_{\rho}}(\hat{x},s;\mu)[h,h].$ (113)

### 21 May 2026 v2

@@ 4.2 Two-phase path-following framework
Recall that $\Phi_{\rho}(x^{(k)},s^{(k)};\mu^{(k)})=2(x^{(k)}-z_{\rho}(x^{(k)},s^{(k)};\mu^{(k)}))$. By (iii) and (iv), when Algorithm 2 terminates, an approximate KKT triple $(x^{(k)},s^{(k)},\lambda^{(k)})$ for the original SCP problem (1) is obtained, satisfying
$\mathcal{A}x^{(k)}=b,\;\mathcal{A}^{*}\lambda^{(k)}+s^{(k)}=c,\;\|\Phi_{\rho}(x^{(k)},s^{(k)};\mu^{(k)})\|\leq 2\kappa\sqrt{\frac{\varepsilon}{\rho}}.$ (66)
4.3 Linear systems in the algorithms
In practical computations, the primal constraint is satisfied at all iterates. Thus, the Newton systems (61) and (40) arising in Algorithms 1 and 2 can be written in the unified form
$\begin{pmatrix}\mathcal{A}&0&0\\ 0&\mathcal{I}_{\mathbb{E}}&\mathcal{A}^{*}\\ \mathcal{H}^{-1}\mathcal{W}&\rho^{-1}\mathcal{H}^{-1}&0\end{pmatrix}\begin{pmatrix}\Delta x\\ \Delta s\\ \Delta\lambda\end{pmatrix}=\begin{pmatrix}0\\ r_{1}\\ r_{2}\end{pmatrix}$, (67)
@@ 4.3 Linear systems in the algorithms
[…]
Directly solving the full Newton system (67) is computationally expensive, particularly for large-scale problems when the dimensions of the cone variables $x$ and $s$ are significantly larger than the number of constraints. To reduce the computational cost, we eliminate $\Delta x$ and $\Delta s$ from (67), which leads to the following Schur-complement system:
$\mathcal{A}\mathcal{W}^{-1}\mathcal{A}^{*}\Delta\lambda$ $=\mathcal{A}\mathcal{W}^{-1}r_{1}-\rho\mathcal{A}\mathcal{W}^{-1}\mathcal{H}r_{2},$ (68a)
$\Delta s$ $=r_{1}-\mathcal{A}^{*}\Delta\lambda,$ (68b)
$\Delta x$ $=\mathcal{W}^{-1}(\mathcal{H}r_{2}-\rho^{-1}\Delta s).$ (68c)
For both IPMs and classical SNMs, the dominant computational cost typically comes from forming and factorizing the Schur complement $\mathcal{A}\mathcal{D}\mathcal{A}^{*}$. In our method, $\mathcal{D}=\mathcal{W}^{-1}$, where $\mathcal{W}$ is an iterate-dependent operator determined by the barrier function in PFSNM. Accordingly, improving the efficiency of forming $\mathcal{A}\mathcal{D}\mathcal{A}^{*}$ is crucial for large-scale computation. To this end, Proposition 3 provides closed-form expressions for $\mathcal{A}\mathcal{D}\mathcal{A}^{*}$ in PFSNM for the three most common symmetric cones.
@@
Let $\mathbb{K}$ be one of the symmetric cones considered below, and let $e$ denote the corresponding Jordan identity element. Then the corresponding Schur complement admits the following explicit representations.
@@
    Let $\mathbb{K}=\mathbb{Q}^{n+1}$. Then, for any $z\in\operatorname{int}(\mathbb{Q}^{n+1})$, $\phi(z)=-\ln\big(z_{0}^{2}-\|\bar{z}\|^{2}\big),$ and
@@
By definition, ${\mathcal{W}}=\dfrac{\mu}{\rho}D^{2}\phi(z_{\rho})$. The results follow directly from the explicit formulas (see (Vieira, 2007, Proposition 2.6.1)) for $D^{2}\phi$.
The importance of explicit Schur-complement formulas has already been noted in the SDP literature. The SNM in Chen and Tseng (2003), based on the smoothing FB function, has a Schur-complement formation cost comparable to that of the most expensive Alizadeh–Haeberly–Overton (AHO) direction in IPMs. If the smoothing CHKS function is used instead, the formation cost becomes lower than that of AHO, but remains higher than that of the Nesterov–Todd (NT) direction. In our earlier work Zhang et al. (2024), we showed that once an explicit formula for the Schur complement is available, the formation cost can be reduced to the same order as that of the NT direction. Compared with Zhang et al. (2024), this paper leverages self-concordant properties to derive this explicit Schur-complement formula in a simpler and more direct way.
For SOCP, the advantage of the explicit Schur complement is even more significant. As shown in case (ii) of Proposition 3, the Schur complement takes the form of a scaled matrix plus two rank-one updates. However, the vector $u:=\mathcal{A}z_{\rho}$ appearing in these rank-one terms is typically dense, which causes the resulting Schur complement to be dense as well. Figure 2 illustrates this structure and shows how these components combine to yield a fully dense matrix. This density poses a major computational challenge in large-scale settings. Even accelerating the evaluation of $\mathcal{D}$ as described in Fukushima et al. (2002) does not solve the problem, because the Schur complement remains dense. The explicit representation in Proposition 3 avoids forming this dense matrix. The terms $\mathcal{A}\mathcal{A}^{*}$ and $\mathcal{A}e$ depend only on the problem data and can be precomputed once. Each subsequent iteration then requires only one matrix-vector product $u=\mathcal{A}z_{\rho}$, along with simple low-rank updates and the scalar $\det(z_{\rho})$. With this structure, one can apply either the product-form Cholesky factorization approach in Alizadeh and Goldfarb (2003) or the expanded sparse representation technique in Zhang et al. (2026). Both approaches exploit the low-rank structure and avoid forming a dense Schur complement.
@@ 5 Complexity analysis
In this section, we establish a polynomial iteration complexity for PFSNM applied to SCP. The analysis consists of two parts. We first derive an iteration complexity for the first phase. Starting from the point produced by this phase, Algorithm 2 then attains an iteration complexity of order $\mathcal{O}(\sqrt{\nu}\ln(1/\varepsilon))$, matching the classical short-step interior-point complexity Vavasis and Ye (1996).
The iteration complexity for the first phase is presented in the following theorem.
@@
Suppose that $w^{(0,0)}\in K(\theta_{\rho})$ and that $t^{(0)}$ in Algorithm 1 satisfies (63). Then Algorithm 1 requires at most
$\mathcal{O}\left(1+\frac{4t^{(0)}}{\kappa}\Theta_{1}\left(\dfrac{\theta_{\rho}(w^{(0,0)};\mu^{(0)})}{\mu^{(0)}}\right)\right)$ (69)
iterations to obtain a starting point $v^{(0)}\in\mathcal{N}(\kappa,\mu^{(0)},\rho)$, where $\Theta_{1}(\cdot)$ is a properly chosen universal positive, continuous, and nondecreasing function on $\mathbb{R}_{+}$.
@@
Since $\mu$ remains fixed in the first phase, we write $\eta_{\rho}(w)$ and $\eta_{t,\rho}(w)$ in place of $\eta_{\rho}(w;\mu^{(0)})$ and $\eta_{t,\rho}(w;\mu^{(0)})$, respectively. Define the merit function associated with $\eta_{t,\rho}$ by
$\delta_{t,\rho}(w):=\|(D^{2}_{ww}\eta_{\rho}(w))^{-1}\nabla_{w}\eta_{t,\rho}(w)\|_{\eta_{t,\rho},w}=\|(D^{2}_{ww}\eta_{\rho}(w))^{-1}\nabla_{w}\eta_{t,\rho}(w)\|_{\eta_{\rho},w}.$
@@
$\delta_{t^{(j+1)},\rho}(w^{(0,j)})-\delta_{t^{(j)},\rho}(w^{(0,j)})$
$=$ $\ \|(D^{2}_{ww}\eta_{\rho}(w^{(0,j)}))^{-1}\nabla_{w}\eta_{t^{(j+1)},\rho}(w^{(0,j)})\|_{\eta_{\rho},w^{(0,j)}}$
$\qquad\qquad\qquad-\|(D^{2}_{ww}\eta_{\rho}(w^{(0,j)}))^{-1}\nabla_{w}\eta_{t^{(j)},\rho}(w^{(0,j)})\|_{\eta_{\rho},w^{(0,j)}}$
$\leq$ $\ \alpha^{(j)}t^{(j)}\|(D^{2}_{ww}\eta_{\rho}(w^{(0,j)}))^{-1}\nabla_{w}\eta_{\rho}(w^{(0,0)})\|_{\eta_{\rho},w^{(0,j)}}$
$\leq$ $\ \frac{\kappa}{4}.$
Here the first inequality follows from the identity for $\nabla_{w}\eta_{t,\rho}(w)$ and from the fact that $\|\cdot\|_{\eta_{\rho},w^{(0,j)}}$ defines a norm. The second inequality follows from (64). Thus,
@@
$\delta_{t^{(j+1)},\rho}(w^{(0,j+1)})\leq\left(\frac{\delta_{t^{(j+1)},\rho}(w^{(0,j)})}{1-\delta_{t^{(j+1)},\rho}(w^{(0,j)})}\right)^{2}\leq\frac{\kappa}{2},$
@@
By applying Nemirovski’s estimate (Nemirovski, 1999, Lemma 8.3(a)) to the standard self-concordant convex-concave function $\frac{\eta_{\rho}(w;\mu^{(0)})}{\mu^{(0)}}$, there exists a universal positive, continuous, and nondecreasing function $\Theta_{1}$ such that
$\|(D^{2}_{ww}\eta_{\rho}(w^{(0,j)}))^{-1}\nabla_{w}\eta_{\rho}(w^{(0,0)})\|_{\eta_{\rho},w^{(0,j)}}\leq\Theta_{1}\left(\frac{\theta_{\rho}(w^{(0,0)};\mu^{(0)})}{\mu^{(0)}}\right).$ (73)
It remains to estimate the number of updates of $t$. If $\alpha^{(j)}=1$, then $t^{(j+1)}=(1-\alpha^{(j)})t^{(j)}=0$. Hence, by (72),
$\delta_{\rho}(w^{(0,j+1)};\mu^{(0)})=\delta_{t^{(j+1)},\rho}(w^{(0,j+1)})\leq\frac{\kappa}{2}.$
Thus Algorithm 1 terminates after at most one additional Newton step.
We next consider the case where $\alpha^{(j)}<1$. Using the definition of $\alpha^{(j)}$ in (64), we have
$\alpha^{(j)}t^{(j)}=\frac{\kappa}{4\left\|(D^{2}_{ww}\eta_{\rho}(w^{(0,j)};\mu^{(0)}))^{-1}\nabla_{w}\eta_{\rho}(w^{(0,0)};\mu^{(0)})\right\|_{\eta_{\rho},w^{(0,j)}}}.$
Therefore,
$t^{(j+1)}$ $=(1-\alpha^{(j)})t^{(j)}$
$=t^{(j)}-\frac{\kappa}{4\left\|(D^{2}_{ww}\eta_{\rho}(w^{(0,j)};\mu^{(0)}))^{-1}\nabla_{w}\eta_{\rho}(w^{(0,0)};\mu^{(0)})\right\|_{\eta_{\rho},w^{(0,j)}}}.$
By (73), this implies
$t^{(j+1)}\leq t^{(j)}-\frac{\kappa}{4\Theta_{1}\left(\frac{\theta_{\rho}(w^{(0,0)};\mu^{(0)})}{\mu^{(0)}}\right)}.$
Consequently, as long as Algorithm 1 has not yet reached a point satisfying
$t^{(j)}\Theta_{1}\left(\frac{\theta_{\rho}(w^{(0,0)};\mu^{(0)})}{\mu^{(0)}}\right)\leq\frac{\kappa}{2},$
the parameter $t^{(j)}$ decreases by at least $\frac{\kappa}{4\Theta_{1}\left(\frac{\theta_{\rho}(w^{(0,0)};\mu^{(0)})}{\mu^{(0)}}\right)}$ at each update. Hence, after at most
$\mathcal{O}\left(\frac{4t^{(0)}}{\kappa}\Theta_{1}\left(\frac{\theta_{\rho}(w^{(0,0)};\mu^{(0)})}{\mu^{(0)}}\right)\right)$
updates, we obtain an index $j$ such that $t^{(j)}\Theta_{1}\left(\frac{\theta_{\rho}(w^{(0,0)};\mu^{(0)})}{\mu^{(0)}}\right)\leq\frac{\kappa}{2}.$ Combining this estimate with (72) and (73), we get
$\delta_{\rho}(w^{(0,j)};\mu^{(0)})$ $=\left\|(D^{2}_{ww}\eta_{\rho}(w^{(0,j)}))^{-1}\nabla_{w}\eta_{\rho}(w^{(0,j)})\right\|_{\eta_{\rho},w^{(0,j)}}$
$\leq\delta_{t^{(j)},\rho}(w^{(0,j)})+t^{(j)}\left\|(D^{2}_{ww}\eta_{\rho}(w^{(0,j)}))^{-1}\nabla_{w}\eta_{\rho}(w^{(0,0)})\right\|_{\eta_{\rho},w^{(0,j)}}$
$\leq\frac{\kappa}{2}+t^{(j)}\Theta_{1}\left(\frac{\theta_{\rho}(w^{(0,0)};\mu^{(0)})}{\mu^{(0)}}\right)\leq\kappa.$
Therefore, Algorithm 1 terminates. The final Newton step computed from (40) then gives a point $v^{(0)}\in\mathcal{N}(\kappa,\mu^{(0)},\rho)$. Thus Algorithm 1 requires at most
$\mathcal{O}\left(1+\frac{4t^{(0)}}{\kappa}\Theta_{1}\left(\dfrac{\theta_{\rho}(w^{(0,0)};\mu^{(0)})}{\mu^{(0)}}\right)\right)$
iterations to obtain a starting point $v^{(0)}\in\mathcal{N}(\kappa,\mu^{(0)},\rho)$.
Having established the iteration complexity for the first phase of PFSNM, we now turn to analyze the iteration complexity of Algorithm 2. As a preliminary step, Theorems 5.2 and 5.3 quantify the effect of updating the smoothing parameter $\mu$ on both $\nabla_{w}\eta_{\rho}(w;\mu)$ and $S_{\eta_{\rho}}(w;\mu)$. For notational simplicity, we omit the arguments $(w;\mu)$ and write, for example, $\nabla_{w}\eta_{\rho}:=\nabla_{w}\eta_{\rho}(w;\mu)$ whenever first- or second-order derivatives of $\eta_{\rho}$ with respect to $w=(\hat{x},s)$ are mentioned. All derivatives with respect to $\mu$ are denoted by a prime.
@@
For any $\mu>0$, $\rho>0$, any direction $h=(h_{\hat{x}},h_{s})\in\hat{\mathbb{E}}\times\mathbb{E}$, and any point $w=(\hat{x},s)\in\hat{\mathbb{E}}\times\mathbb{E}$,
$\left|\langle h,\nabla_{w}\eta_{\rho}^{\prime}(w;\mu)\rangle\right|\leq\sqrt{\dfrac{2\nu}{\mu}}\sqrt{S_{\eta_{\rho}}(w;\mu)[h,h]}.$ (74)
@@
$\nabla_{w}\eta_{\rho}^{\prime}(w;\mu)=\begin{pmatrix}-\mathcal{B}^{*}y^{\prime}_{\rho}\\ z^{\prime}_{\rho}\end{pmatrix}=\begin{pmatrix}\mathcal{B}^{*}\mathcal{H}^{-1}\nabla\phi(z_{\rho})\\ -\rho^{-1}\mathcal{H}^{-1}\nabla\phi(z_{\rho})\end{pmatrix}$.
@@
$\langle\nabla_{w}\eta_{\rho}^{\prime}(w;\mu),h\rangle=\langle\mathcal{B}h_{\hat{x}},\mathcal{H}^{-1}\nabla\phi(z_{\rho})\rangle-\rho^{-1}\langle h_{s},\mathcal{H}^{-1}\nabla\phi(z_{\rho})\rangle.$
@@
$\left|\langle\nabla_{w}\eta^{\prime}_{\rho},h\rangle\right|$ $\leq\left|\langle h_{x},\mathcal{H}^{-1}\nabla\phi(z_{\rho})\rangle\right|+\rho^{-1}\left|\langle h_{s},\mathcal{H}^{-1}\nabla\phi(z_{\rho})\rangle\right|$ (75)
@@
$\qquad\qquad\qquad+\rho^{-1}\sqrt{\langle h_{s},\mathcal{H}^{-1}h_{s}\rangle}\sqrt{\langle\nabla\phi(z_{\rho}),\mathcal{H}^{-1}\nabla\phi(z_{\rho})\rangle}.$
Since $\mathcal{H}^{-1}=(\mathcal{I}_{\mathbb{E}}+\mathcal{W})^{-1}$, we have
$\mathcal{W}^{-1}\mathcal{H}^{-1}\prec\mathcal{W}^{-1},\quad\mathcal{H}^{-1}\prec\mathcal{W}^{-1},$
which together with $\mathcal{W}=\frac{\mu}{\rho}D^{2}\phi(z_{\rho})$ and (10) implies
$\sqrt{\langle\nabla\phi(z_{\rho}),\mathcal{W}^{-1}\mathcal{H}^{-1}\nabla\phi(z_{\rho})\rangle}$ $\leq\sqrt{\frac{\rho}{\mu}\langle\nabla\phi(z_{\rho}),(D^{2}\phi(z_{\rho}))^{-1}\nabla\phi(z_{\rho})\rangle}=\sqrt{\frac{\rho\nu}{\mu}},$ (76)
$\sqrt{\langle\nabla\phi(z_{\rho}),\mathcal{H}^{-1}\nabla\phi(z_{\rho})\rangle}$ $\leq\sqrt{\frac{\rho}{\mu}\langle\nabla\phi(z_{\rho}),(D^{2}\phi(z_{\rho}))^{-1}\nabla\phi(z_{\rho})\rangle}=\sqrt{\frac{\rho\nu}{\mu}}.$
Combining (75), (76) and Corollary 1 yields
$|\langle\nabla_{w}\eta^{\prime}_{\rho},h\rangle|$ $\leq\sqrt{\dfrac{\nu}{\mu}}\left(\sqrt{\langle h_{\hat{x}},D_{\hat{x}\hat{x}}^{2}\eta_{\rho}h_{\hat{x}}\rangle}+\sqrt{-\langle h_{s},D_{ss}^{2}\eta_{\rho}h_{s}\rangle}\right)$
@@
This completes the proof.
@@
For any $\mu>0$, $\rho>0$, any direction $h=(h_{\hat{x}},h_{s})\in\hat{\mathbb{E}}\times\mathbb{E}$, and any point $w=(\hat{x},s)\in\hat{\mathbb{E}}\times\mathbb{E}$,
$\left|S_{\eta_{\rho}}^{\prime}(w;\mu)[h,h]\right|\leq\frac{1+{2\sqrt{\nu}}}{\mu}S_{\eta_{\rho}}(w;\mu)[h,h].$ (77)